OpenStudy (anonymous):

x^2+7x+10=y solve quadratic formula

OpenStudy (anonymous):

what does a=? b=? c=? In this problem?

OpenStudy (anonymous):

it doesn't say

OpenStudy (anonymous):

I know it doesnt you have to figure it out by looking in the problem.

OpenStudy (anonymous):

oh. c=-10? and b=7x?

OpenStudy (anonymous):

c is going to be positive so you only have to change the sign. b is the number next x in 7x we only want the number not the letter. a is the number next to x^2 hint for this one: it's not 2 because thats the exponent we want the number left of x since in this case we have no number what will it be?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yeah you are correct a=1 b=7 c=10

OpenStudy (anonymous):

Now plug those numbers in for \begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}

OpenStudy (anonymous):

\[\pm \] this signs means you have to do it two times. Once when that sign becomes + and the other is when you replace it with + that sign means plus and minus

OpenStudy (anonymous):

\[-7\pm \sqrt{?}\]

OpenStudy (anonymous):

9/2